OpenStudy (yanni):

Slopes of parallel and perpendicular lines: Problem type 1 Consider the line . (a) What is the slope of a line parallel to this line? (b) What is the slope of a line perpendicular to this line?

OpenStudy (yanni):

the line is 8x-4y=7

OpenStudy (anonymous):

when a line is parallel to a line then Slop1=Slop2 When A line is Perpendicular to a line then slop1*slop2=-1

OpenStudy (anonymous):

Do u know what is the slope of 8x-4y=7 ?

OpenStudy (yanni):

y=2x-7/4?

OpenStudy (anonymous):

Thats the 'slope-intercept' formula y=mx+b where m is the slope according to that The slope here is 2 , Agree ?

OpenStudy (yanni):

ya

OpenStudy (anonymous):

OK ,Point (A) when a line is parallel to a line then Slope1=Slope2 U have Slope 1 ,can u get slope 2 ?

OpenStudy (anonymous):

?

OpenStudy (yanni):

a=2 and b=-1/2

OpenStudy (anonymous):

Correct! :) Good Job .